# Online Episodic Memory Visual Query Localization with Egocentric Streaming Object Memory

### v1

@@ Abstract
[…]
denotes equal contribution. denotes equal supervision. Correspondence can be sent to: matteo.dunnhofer@uniud.it
@@ 1 Introduction
Day-to-day experiences often witness lapses in human memory. We may misplace objects within our homes (like searching for a passport), lose track of completed tasks (wondering if salt was added), forget the locations of past activities (where were tickets purchased?), and overlook the status of objects in our surroundings (did I leave the garage door open?). Wearable devices observing human activities from an egocentric perspective have the potential to assist the recall of such episodic memories [43] in a wide range of scenarios by continuously keeping track of observed objects and events to alleviate cognitive overload and support the execution of daily tasks [33].
@@
Since daily activities revolve around interactions with objects serving specific functions, such as tools and appliances, it is natural to design episodic memory retrieval systems around the ability to correctly recall the occurrence of specific object instances. Towards this direction, the Visual Queries (VQ) task has been introduced with the Episodic Memory (EM) benchmark of the Ego4D dataset [18]. The VQ task is formulated as follows: given an image patch of an object query and a long egocentric video, identify a response track, the precise spatio-temporal location of the last appearance of the object. The interest of the community in EM tasks has been fostered by the regular organization of international challenges, which recently spurred a number of technical solutions [25, 13, 6, 31, 46] and research works [47, 35, 22, 34, 29]. While demonstrating the feasibility of addressing the problem, all previously presented solutions operate in an offline manner. Specifically, given a visual query at frame $t$, they search within the same video, scanning it from current frame $t$ backward up to the moment in which the object is found. This translates into algorithms that, for each new query object, necessitate storing and reprocessing all the frames previously observed in the video stream. We argue that this approach is unfeasible in practical applications, especially considering a scenario in which an always-on wearable device processes several hours of video per day. [^2] For instance, a 24-hour 8-bit RGB video with 1408x1408 resolution captured at 20FPS [40] can take up to 3.43 TB of memory uncompressed. Additionally, processing different queries within the same video entails executing the algorithm several times. In essence, we argue that, despite the potential of current algorithms, an offline query object search is not practically applicable due to its significant computational and storage requirements.
@@
In sharp contrast with this burdensome and offline scenario, we advocate that EM systems should be designed to work in an online fashion, with limited budgets for computation and storage reflecting the need to ultimately deploy the algorithms on wearable devices. Note that an online formulation of the episodic memory task brings several challenges, beyond the need for efficient systems. Specifically: 1) An online system has to identify objects which could be potentially queried in the future without any prior knowledge of their identity, in contrast with offline systems which begin the object search only after the object query is known; 2) The online system should be able to keep track of the different occurrences of a given object instance and form appropriate response tracks in advance, in contrast with offline systems which initiate a separate search for each query; 3) The online system should be able to associate each of the potential response tracks with appropriate representations to enable future retrieval, in contrast with an offline system, which can directly compare the query patch with object candidates during the search process.
To address this challenge, in this paper, we propose ESOM - Egocentric Stream Object Memory, a novel architecture which processes egocentric video frames in an online fashion and keeps a low-dimensional, interpretable dynamic memory of all objects observed in past video (see Figure 1). This is achieved thanks to three main components: 1) an object discovery module which detects objects of interest in each input frame; 2) an online, long-term visual object tracker which takes over the discovered objects to track them in the video; 3) a dynamic object memory which stores past observations in the form of object tracks with associated visual representations. The object memory is maintained by an object memory population process which keeps the memory updated by running the object discovery and visual object tracker components at any frame, and can be queried by a query retrieval and localization process which allows to efficiently search within the memory when a new query is made and return the appropriate response track. Experiments on the established Ego4D episodic memory retrieval benchmark [18] show the effectiveness of the proposed ESOM architecture, which outperforms offline episodic memory retrieval methods when oracular object discovery and tracking components are considered, achieving a success rate of $81.92\%$ vs $55.89\%$ of the best offline method. We also show that results in real-world non-oracular settings are hindered by the limited performance of current detection and tracking algorithms, which suggests that further research efforts should be devoted to the development of these components, for which this work offers a principled benchmark grounded in the real-world downstream task of episodic memory retrieval.
@@
The contributions of this work are as follows: 1) we are the first to introduce the novel task of online episodic memory retrieval; 2) we propose ESOM, a novel architecture based on continuous object discovery, tracking, memory update, and query, which is shown to be in principle competitive with offline methods; 3) we offer a benchmark based on Ego4D [18] of online episodic memory retrieval in general and object detection/tracking in egocentric vision.
@@ 2 Related Work
The problem of Visual Query Localization was initially formalized in [18], along with a baseline cascade solution that first searches for the query patch in every frame by means of a Siamese R-CNN [44], and then tracks it forward and backward in time with a short-term tracker [4] to obtain the temporal extent of the response track. Based on this formulation, a variety of methodologies explored different key factors such as balancing positive and negative examples [46], addressing domain and task biases [47], improving feature extractors [6, 31] , and developing end-to-end single-stage frameworks  [22] to improve performance and inference speed. Despite the progress in this area, all the aforementioned works are based on the offline assumption. In this paper, we formulate a novel Online Episodic Memory Visual Query Localization task (OEM-VQL) and offer a first approach and benchmark to tackle the task.
@@
Few previous works investigated the use of object memory to support cognition, manipulation, or video understanding in robotics [21], to support object permanence in manipulation tasks, and in egocentric vision [32], to mimic human spatial cognition keeping track of object instances in 3D when they are either in-sight or out-of-sight. Recent work [11, 17], in particular, explored the use of an object memory to address video understanding tasks such as generic question answering  [11] and comprehension of object-based activities in long egocentric videos [17]. Differently from the aforementioned works, our methods is the first to propose and tackle online episodic memory with an object memory, which raises significant challenges, such as dealing with egocentric video, accounting for all objects appearing in the scene, and accurately track object instances in the video.
@@
Some previous works tackled the design of algorithms and systems which to support online processing, a fundamental requirement for real-world applications on wearable devices [8, 50, 1, 20, 38, 41, 7, 15, 14, 37, 16], in contrast to a whole line of egocentric vision tasks such as action recognition [5, 12, 3] and temporal action detection [49, 27] which assume that the whole video is available at inference time. Episodic memory retrieval tasks, as initially formulated in [18] are based on the “offline” assumption. Limited amount of works considered an alternative online formulation which resembles an online setting [2], but a systematic investigation on the topic is still missing. We address the problem of online episodic memory retrieval targeting the VQL task. The developed system achieves online processing abilities by making use of visual object tracking and a compact queryable memory module to store past experiences.
3 Online Episodic Memory - Visual Queries
This paper addresses the challenge of locating an image patch of a query object instance within an online video stream that cannot be stored. We refer to this problem as the Online Episodic Memory - Visual Query Localization (OEM-VQL) task. More formally, given a streaming egocentric video $\mathcal{V}_{ego}=\big{\{}\mathbf{F}_{t}\big{\}}$ of RGB frames $\mathbf{F}_{t}$ indexed by time $t$, and a query RGB image patch $\mathbf{Q}$ containing the visual appearance of the query object, the goal is to return a response track:
$\mathbf{r}=\big{\{}\mathbf{b}_{s},\mathbf{b}_{s+1},\dots,\mathbf{b}_{e-1},\mathbf{b}_{e}\big{\}},$ (1)
i.e., an ordered sequence of bounding-boxes $\mathbf{b}_{j},s\leq j\leq e\leq t$ which determines the spatio-temporal position of the last appearance of the object in $\mathcal{V}_{ego}$. $s$ and $e$ are the starting and ending frames of the response track and each box $\mathbf{b}_{j}$ corresponds to frame $\mathbf{F}_{j}$. As in [18], the goal is to find $\mathbf{r}$ given $\mathbf{Q}$, however, critically, our formulation does not grant access to the complete video $\mathcal{V}_{ego}$. Rather, we allow methods to exploit just the visual information in $\mathbf{F}_{t}$ plus compact object memories computed at previous frames $\mathbf{F}_{\hat{t}}$ with $\hat{t}<t$. Figure 1 provides a visual illustration of our problem formulation in contrast to the one adopted in [18].
@@ 4 Egocentric Streaming Object Memory
To tackle OEM-VQL, we propose the Egocentric Streaming Object Memory (ESOM) architecture. ESOM is based on the object memory $\mathcal{M}_{ego}$, a dynamic data structure that memorizes relevant object instances appearing in $\mathcal{V}_{ego}$. The data structure is populated and queried through two processes: (i) the object memory population algorithm that builds such data structure online and continuously as the video $\mathcal{V}_{ego}$ is being captured; (ii) the query retrieval and localization algorithm that retrieves the response track $\mathbf{r}$ related to the visual query $\mathbf{Q}$ from the constructed data structure, whenever needed by the camera wearer. The continuous interaction between these components allows to dynamically create, update, and query the memory of previously observed objects to support online episodic memory retrieval without the need of storing all the frames of $\mathcal{V}_{ego}$ (see Figure 1).
@@ 4.1 Object Memory
The core component of the ESOM architecture is the object memory $\mathcal{M}_{ego}$, a dynamic data structure which maintains spatio-temporal information of object instances relevant to the camera-wearer and is updated at every frame $\mathbf{F}_{t}$.
@@
of compact, high-level representations $\mathbf{O}_{i}$ of the spatio-temporal locations of objects in the frames $\mathbf{F}_{t}$ of $\mathcal{V}_{ego}$. The indexes $1\leq i\leq N$ denote object instance unique IDs. While the representation of $\mathbf{O}_{i}$ can be arbitrarily abstract, in practice, we set $\mathbf{O}_{i}$ to be a list $\mathbf{O}_{i}=\big{\{}\mathbf{o}_{i,t}\big{\}}$ of records temporally indexed by $t$, with each record having the following form
$\mathbf{o}_{i,t}=\big{[}\mathbf{b}_{i,t},\phi_{i,t}\big{]}$ (3)
where $\mathbf{b}_{i,t}=[b_{i,t}^{x},b_{i,t}^{y},b_{i,t}^{w},b_{i,t}^{h}]$ denotes the object bounding-box, and $\phi_{i,t}$ is a representation of the object’s appearance (e.g., in the form of RGB patches or extracted features). If $\mathbf{o}_{i,t}$ is not visible in $\mathbf{F}_{t}$, then $\mathbf{o}_{i,t}=\emptyset$. This formulation for the object memory provides an indexed, lightweight, compact, human-readable and easily queryable representation of objects observed in $\mathcal{V}_{ego}$ up to the current frame $\mathbf{F}_{t}$. Figure 2 visualizes the representation of $\mathcal{M}_{ego}$ with respect to the visual information of $\mathcal{V}_{ego}$.
@@
The .$\texttt{write}\texttt{}$ operation allows to store new object instances or update existing ones. As before, this operation can be called with different input parameters. The syntax
$\mathcal{M}_{ego}.\texttt{write}\texttt{(}\big{[}\mathbf{b}_{k,t},\phi_{k,t}\big{]}\texttt{)}$ (5)
stores with ID $k$ (not yet present in $\mathcal{M}_{ego})$ a new object instance $\mathbf{O}_{k}$ by inserting the bounding-box $\mathbf{b}_{k,t}$ and the representation $\phi_{k,t}$. The write operation can also be called as
$\mathcal{M}_{ego}.\texttt{write}\texttt{(}\big{[}\mathbf{b}_{i,t},\phi_{i,t}\big{]}\texttt{)}$ (6)
to add bounding-box $\mathbf{b}_{i,t}$ and visual representation $\phi_{i,t}$ to object instance $i$ at time $t$ (i.e., we update the representation of $\mathbf{O}_{i}$). To create a consistent, yet compact, memory, the .$\texttt{write}\texttt{()}$) operation implements a filter to select which pieces of object information to be saved in $\mathcal{M}_{ego}$. Information not useful to visual query retrieval should be excluded from $\mathcal{M}_{ego}$. For instance, $\phi_{i,t}$ might not contain sufficiently distinctive features for accurate instance recognition during retrieval. To avoid these situations, .$\texttt{write}\texttt{(}\big{[}\mathbf{b}_{i,t},\phi_{i,t}\big{]}\texttt{)}$ selectively saves only high-quality visual representations by replacing the previously memorized one, except for the first effective representation of the current tracklet. This selection is guided by a binary classification function that assigns a probability score to each $\phi_{i,t}$, indicating whether it is of sufficient quality for effective query retrieval (see Supp. material for details). In case of a negative decision, the visual representation is set to $\phi_{i,t}=\emptyset$ and only the bounding-box is memorized. In this way, we retain the representation of the object at the moment of discovery, as well as two representations from the most recent response track that are as distant as possible from each other and considered effective by the classifier.
Figure 3: Memory population by object tracking and discovery. Between consecutive frames $F_{t-1}$ and $F_{t}$, the Object Tracking (OT) module of the OMP algorithm reads the object representations $\mathbf{o}_{i,t-1}$ (if stored in $\mathcal{M}_{ego}$), updates objects’ positions to $\mathbf{b}_{i,t}$, extracts representations $\phi_{i,t}$, and writes this information into $\mathcal{M}_{ego}$. In parallel, new objects to store in $\mathcal{M}_{ego}$ are detected by the Object Discovery (OD) module in frame $F_{t}$. This provides bounding-boxes of relevant object instances $\mathbf{b}_{k,t}$, extracts representations $\phi_{k,t}$ from them, and writes them to $\mathcal{M}_{ego}$.

### v2

@@ Abstract
Episodic memory retrieval enables wearable cameras to recall objects or events previously observed in video. However, existing formulations assume an “offline” setting with full video access at query time, limiting their applicability in real-world scenarios with power and storage-constrained wearable devices. Towards more application-ready episodic memory systems, we introduce Online Visual Query 2D (OVQ2D), a task where models process video streams online, observing each frame only once, and retrieve object localizations using a compact memory instead of full video history. We address OVQ2D with ESOM (Egocentric Streaming Object Memory), a novel framework integrating an object discovery module, an object tracking module, and a memory module that find, track, and store spatio-temporal object information for efficient querying. Experiments on Ego4D demonstrate ESOM’s superiority over other online approaches, though OVQ2D remains challenging, with top performance at only  4% success. ESOM’s accuracy increases markedly with perfect object tracking (31.91%), discovery (40.55%), or both (81.92%), underscoring the need of applied research on these components.
denotes equal contribution. denotes equal supervision.
@@ 1 Introduction
Day-to-day experiences often witness lapses in human memory. We may misplace objects within our homes (like searching for a passport), lose track of completed tasks (wondering if salt was added), forget the locations of past activities (where were tickets purchased?), and overlook the status of objects in our surroundings (did I leave the garage door open?). Wearable cameras observing human activities from an egocentric perspective have the potential to assist the recall of such episodic memories [47] in a wide range of scenarios by continuously keeping track of observed objects and events to alleviate cognitive overload and support the execution of daily tasks [36].
@@
Towards this direction, the Visual Queries 2D (VQ2D) task has been introduced with the Episodic Memory (EM) benchmark of the Ego4D dataset [22]. VQ2D is formulated as follows: given an image patch of an object query and a long egocentric video, identify a response track, the precise spatio-temporal location of the last appearance of the object in the video. VQ2D has fostered the development of algorithms for episodic memory retrieval in recent years [29, 17, 7, 35, 50, 51, 39, 26, 38, 33]. We observe that the current formulation of VQ2D provides for solutions to operate in an offline manner. Specifically, given a visual query at frame $t$, methods search within $\mathcal{V}_{ego}$ scanning it from current frame $t$ backward up to the moment in which the object is found. Arguably, this approach is limited in practical application scenarios in which an always-on wearable device processes several hours of video per day. In fact, offline VQ2D requires the burdensome storage of all past videoFor instance, a 24-hour 8-bit RGB video with 1408x1408 resolution captured at 20FPS [44] can take up to 3.43 TB of uncompressed memory. and entails significant computational overhead when retrieving multiple queries, requiring the scan of the same video several times. We advocate that EM systems should be able to work in an online fashion, with limited budgets for computation and storage, reflecting the need to ultimately deploy algorithms on wearable devices.
@@
We hence propose a new task formulation for EM, Online Visual Query 2D (OVQ2D), in which models are required to work online, processing video frames only once, storing relevant information in a compact memory, and relying on such memory to answer queries whenever they are made by the user. Similarly to offline VQ2D, methods need to return a series of bounding boxes and corresponding past frames where the queried object appears. This enables clear spatio-temporal localization within its visual context, forming an interpretable episode of its last known location. To assess practical applicability, we compare memory size and computation time, besides mere retrieval performance. Note that this online formulation brings several challenges, beyond the need for efficient systems. Specifically: 1) online systems have to identify objects which could be potentially queried in the future without any prior knowledge of their identity, in contrast to offline systems which begin the object search only after the query is known; 2) online systems should be able to keep track of the different occurrences of a given object instance and store appropriate response tracks and related frames in advance, in contrast to offline systems which initiate a separate search for each query; 3) Online systems should be able to associate each of the potential response tracks with appropriate representations to enable future retrieval, in contrast to offline systems, which can directly compare the query patch with object candidates during the search process.
To tackle OVQ2D, we propose ESOM (Egocentric Stream Object Memory), a principled architecture built on state-of-the-art object detection and tracking, leveraging decades of progress in computer vision. This formulation adopts an application-driven focus, prioritizing practical deployment and benchmarking over algorithmic novelty. ESOM processes egocentric video frames online, maintaining a compact, interpretable, and dynamic memory of past objects (see Figure 1). This is achieved with three main components: 1) an object discovery module detecting objects of interest in each input frame; 2) an online, long-term visual object tracker tracking the discovered objects in the video; 3) a dynamic object memory which stores past observations in the form of object tracks with associated visual representations. The object memory is built through an object memory population process that coordinates the object discovery and visual tracking modules. It can then be queried by a query retrieval and localization process, which efficiently searches the memory and returns the corresponding response track in response to a user query.
Experiments on Ego4D [22] demonstrate the effectiveness of ESOM over other online methods. Results also show that OVQ2D is challenging, with the best method achieving a success rate of only $\sim$4%. This modest performance is mainly due the limitations of current object detection and tracking on real-world egocentric videos. When using oracle detection or tracking, ESOM’s success rate rises to 31.91% and 40.55%, respectively, and reaches 81.92% with both. These results also shed light on the limited performance of current object detection and tracking algorithms when employed in real-world applications, suggesting that more efforts are needed to improve their practical applicability. Unlike standard evaluations that assess these components in isolation, our study offers a competitive benchmark that can be used to assess progress and applicability in real-world downstream tasks.
@@
Overall, the contributions of this work are as follows: 1) we introduce OVQ2D, an online formulation for episodic memory retrieval; 2) we propose ESOM, a novel architecture based on continuous object discovery, tracking, memory update, and query, which is shown to outperform offline methods under competitive memory budgets and brings benefits in terms of runtime; 3) we offer a benchmark based on Ego4D [22] for OVQ2D, which we also show as a valuable resource to evaluate real-world performance of object detection and tracking algorithms.
@@ 2 Related Work
The task of Visual Query Localization was initially formalized in [22], along with a baseline cascade solution that first searches for the query patch in every frame by means of a Siamese R-CNN [48], and then tracks it forward and backward in time with a short-term tracker [5] to obtain the temporal extent of the response track. Based on this formulation, a variety of methodologies explored different key factors such as balancing positive and negative examples [50], addressing domain and task biases [51], improving feature extractors [7, 35], and developing end-to-end single-stage frameworks  [26] to improve performance and inference speed. Despite the progress, all the aforementioned works are based on the offline assumption. In this paper, we formulate a novel Online Visual Query 2D (OVQ2D) task and offer a first approach and benchmark to tackle the problem.
@@
Object memory representations have been investigated in robotics [25] to support object permanence in manipulation tasks, and in egocentric vision [37] to mimic human spatial cognition in keeping track of object instances when they are out-of-sight. Recent work [15, 21], in particular, explored the use of an object memory to address video understanding tasks such as generic question answering  [15] and comprehension of object-based activities in long egocentric videos [21]. Unlike the aforementioned works, our method is the first to propose and tackle online episodic memory with an object memory, which raises significant challenges due to the need to account for all objects appearing in the scene and accurately track object instances in the video.
@@
Previous work tackled the design of algorithms and systems to support online processing, a fundamental requirement for real-world applications on wearable devices [10, 55, 2, 24, 42, 45, 9, 19, 18, 41, 20], in contrast to a whole line of egocentric vision tasks such as action recognition [6, 16, 4], and temporal action detection [53, 31] which assume that the whole video is available at inference time. Episodic memory retrieval tasks, as initially formulated in [22] are based on the “offline” assumption. Limited amount of work considered alternative formulations resembling an online setting [3], and a systematic investigation on the topic is still missing. We address the problem of online episodic memory retrieval targeting the VQL task. Our system achieves online processing abilities by using visual object tracking and a compact queryable memory module to store past experiences.
3 Online Visual Query 2D Localization
We define Online Visual Query 2D (OVQ2D) as follows. Given a streaming egocentric video $\mathcal{V}_{ego}=\big{\{}\mathbf{F}_{t}\big{\}}$ of RGB frames $\mathbf{F}_{t}$ indexed by time $t$, and a query image patch $\mathbf{Q}$ representing the visual appearance of the query object, the goal is to retrieve a visual response track:
$\mathbf{r}=\big{\{}(\mathbf{F}_{s},\mathbf{b}_{s}),(\mathbf{F}_{s+1},\mathbf{b}_{s+1}),\dots,(\mathbf{F}_{e},\mathbf{b}_{e})\big{\}},$ (1)
which is an ordered sequence of (frame, bounding box) pairs $(\mathbf{F}_{j},\mathbf{b}_{j})$, where $s\leq j\leq e\leq t$. This track determines the spatio-temporal position of the object’s last appearance in $\mathcal{V}_{ego}$, with $s$ and $e$ denoting starting and ending frames, and each box $\mathbf{b}_{j}$ corresponding to frame $\mathbf{F}_{j}$. This formulation allows for the visualization of $\mathbf{Q}$’s spatio-temporal localization within its surrounding context, generating a detailed video snippet that helps users enhance their natural episodic memory. As in [22], the goal is to find $\mathbf{r}$ given $\mathbf{Q}$, however, critically, our formulation requires methods to process $\mathcal{V}_{ego}$ in a streaming fashion, observing frames $\mathbf{F}_{t}$ only once. To support future queries, methods can store information in a compact memory. Figure 1 provides an illustration of our problem formulation.
@@ 4 Egocentric Streaming Object Memory
To tackle OVQ2D, we propose the Egocentric Streaming Object Memory (ESOM) architecture. ESOM is based on the object memory $\mathcal{M}_{ego}$, a dynamic data structure that memorizes relevant object instances appearing in $\mathcal{V}_{ego}$. The data structure is populated and queried with two processes: (i) the object memory population algorithm that builds $\mathcal{M}_{ego}$ online and continuously as $\mathcal{V}_{ego}$ is being captured; (ii) the query retrieval and localization algorithm that retrieves the visual response track $\mathbf{r}$ related to the query $\mathbf{Q}$ from $\mathcal{M}_{ego}$, whenever needed by the camera wearer. The continuous interaction between these components allows to dynamically create, update, and query the memory of previously observed objects to support online episodic memory retrieval without the need of storing all the frames of $\mathcal{V}_{ego}$.
@@ 4.1 Object Memory
of representations $\mathbf{O}_{i}$ of the spatio-temporal locations of objects in the frames $\mathbf{F}_{t}$ of $\mathcal{V}_{ego}$. The indexes $1\leq i\leq N$ denote object instance unique IDs. While the representation of $\mathbf{O}_{i}$ can be arbitrarily abstract, in practice, we set $\mathbf{O}_{i}$ to be a list $\mathbf{O}_{i}=\big{\{}\mathbf{o}_{i,t}\big{\}}$ of records temporally indexed by $t$, with each record having the following form
$\mathbf{o}_{i,t}=\big{[}t,\mathbf{b}_{i,t},\mathbf{F}_{t},c_{i,t}\big{]}$ (3)
where $\mathbf{b}_{i,t}=[b_{i,t}^{x},b_{i,t}^{y},b_{i,t}^{w},b_{i,t}^{h}]$ denotes the object bounding box, $\mathbf{F}_{t}$ is the associated RGB frame at time $t$, and $c_{i,t}$ is a binary label indicating whether the visual content within $\mathbf{b}_{i,t}$ is relevant (i.e., the appearance is distinctive) for the retrieval process. If $\mathbf{o}_{i,t}$ is not visible in $\mathbf{F}_{t}$, then $\mathbf{o}_{i,t}=\emptyset$. This formulation of the object memory provides an indexed, lightweight, compact, human-readable, and easily queryable representation of objects observed in $\mathcal{V}_{ego}$ up to the current frame $\mathbf{F}_{t}$. Figure 2 visualizes the representation of $\mathcal{M}_{ego}$ with respect to the information of $\mathcal{V}_{ego}$.
@@
The .$\texttt{write}\texttt{()}$) operation allows to store new object instances or update existing ones. As before, this operation can be called with different input parameters. The syntax
$\mathcal{M}_{ego}.\texttt{write}\texttt{(}\big{[}t,\mathbf{b}_{k,t},\mathbf{F}_{t},c_{k,t}\big{]}\texttt{)}$ (5)
stores with ID $k$ (not yet present in $\mathcal{M}_{ego})$ a new object instance $\mathbf{O}_{k}$ by inserting timestamp $t$, the bounding box $\mathbf{b}_{k,t}$, the associated frame $\mathbf{F}_{t}$, if it is not already in $\mathcal{M}_{ego}$, and the label $c_{k,t}$. The write operation can also be called as
$\mathcal{M}_{ego}.\texttt{write}\texttt{(}\big{[}t,\mathbf{b}_{i,t},\mathbf{F}_{t},c_{i,t}\big{]}\texttt{)}$ (6)
to add $t$, $\mathbf{b}_{i,t}$, $\mathbf{F}_{t}$, and $c_{i,t}$ to object instance $\mathbf{O}_{i}$ at time $t$ (i.e., we update the representation of $\mathbf{O}_{i}$). If $\mathbf{F}_{t}$ is already stored in $\mathcal{M}_{ego}$, no duplicate is stored. When .$\texttt{write}\texttt{()}$) updates an object, it checks for a break from the previous timestamp $\hat{t}$ (i.e., $t-\hat{t}>1$). If occurs, frames $\mathbf{F}_{\bar{t}},\bar{t}\leq\hat{t}$ linked to the prior sequence of contiguous boxes are deleted, unless they belong to another memorized object.
[…]